技術経営特論 2 単位 (選択)

Management of Technology  山中英生 · 教授/ 知的力学システム工学専攻建設創造システム工学コース社会システム工学講座

森本恵美 · 助教/ 知的力学システム工学専攻建設創造システム工学コース社会システム工学講座，非常勤講師

【授業目的】自動車産業の生産マネジメントについて，「研究開発」，「生産管理」，「調達」，「品質管理」，「マーケティング」，そして「異文化経営」などの視点から多面的に理解する．

【授業概要】「生産マネジメント」に関する総合的かつ専門的な知識の提供を行う．今回は，特に「自動車産業」，その中でも「本田技研工業 (以下ホンダ)」の生産マネジメントをケースに取り上げ，「マネジメント」における「技術」の位置づけと，その重要性を理解するための授業とする．授業は講義，ケース報告，およびケーススタディから構成され，講義は出口，ケース報告はゲストスピーカー (ホンダ OB 講師)，ケーススタディは出口とゲストスピーカーの連携のもと実施する．なお，この授業は「ニュービジネス概論」，および「ニュービジネス特論」の応用科目として位置づけられており，上記科目と同様に徳島県からの経済的支援を受けて開講される授業科目である．

【授業形式】 [形態] 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】「ニュービジネス概論」もしくは「ニュービジネス概論」を履修し

ていること．【履修上の注意】受講者は毎回必ず十分な予習を行った上で授業に臨むこと．また，

最低でも毎回一人一回は何らかの発言を行うこと．【到達目標】

1. 「マネジメント」の基礎知識を習得する．

2. 「マネジメント」における「技術」の位置づけと，その重要性を理解する．

3. ケース · スタディやビジネス · ゲームを通じて，現場において絶えず発生す

るさまざまな課題の解決方法を学ぶ．【授業計画】

1. ガイダンス

2. 講義「技術経営とは」

3. ケース報告「ホンダの研究開発」

4. ケーススタディ「新車開発指令」

5. 講義「生産 · 調達活動」

6. 講義「品質管理」 7. ケース報告「ホンダの生産活動」，「ホンダの調達活動」

8. ケーススタディ「海外工場を計画せよ」

9. ケース報告「ホンダの品質管理」

10. ケーススタディ「品質不良を再発させるな」

11. 講義「マーケティング戦略」

12. 講義「異文化経営」

13. ケース報告「ホンダの商品企画 ∼ 二輪車を中心に ∼ 」

14. ケース報告「ホンダ国際化の軌跡」

15. ケース報告「ホンダの人事管理」

16. まとめ，および質疑応答

【成績評価基準】授業中の発言 (20%)，および中間 · 期末レポート (合計 80%) で評

価する．【教科書】特に使用しない．毎回資料 · レジュメ等を配付する．【参考書】授業中に数冊紹介する．【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216616 【対象学生】博士前期課程に在籍する全学生のうち，「ニュービジネス概論」，もしくは「ニュービジネス特論」の単位を修得ずみの者を対象学生とする．また，集中的な学習を行うために，受講者数の制限を行う (20∼ 25 名を上限とする予定)．受講者多数の場合は何らかの手段で選抜を行う．

【連絡先】

⇒ 山中 (A410, 088-656-7350, yamanaka@ce.tokushima-u.ac.jp) ���� (オフィスアワー: オフィスアワー:年度ごとに学科の掲示板を参照のこと) ⇒ 森本 (A417, 088-656-7619, emi morimoto@ce.tokushima-u.ac.jp) ���� (オ

フィスアワー: 月 ∼ 木 (10 時 ∼ 16 時)) 【備考】この授業では，「研究開発」，「生産管理」，「調達」，「品質管理」，「マーケ

ティング」，そして「国際人事管理」などの職務に携わってこられた 7 名のホンダ OB を講師として招へいし，ケース報告を行う．こうした試みは全国的に見てもきわめて珍しく，貴重な授業であることを付言しておく．なお，受講希望者は，事前に開催予定のガイダンスに必ず出席すること．日程は別途指示する．

Management of Technology 2 units (selection)

Hideo Yamanaka · Professor / Planning and Design Systems Engineering for Infrastructures, Civil and Environmental Engineering, Intelligent Structures and Mechanics Systems Engineering

Emi Morimoto · Assistant Professor / Planning and Design Systems Engineering for Infrastructures, Civil and Environmental Engineering, Intelligent Structures and Mechanics Systems Engineering

, Part-time Lecturer

Target⟩ About the production management of the automobile industry, it is understood variously from the point of view such as ”research and development”, ”production management”, ”purchasing” and ”quality control” ”marketing” and ”cross-cultural management”.

Outline⟩ It learns about the production management. This time, ”the automobile industry” specially takes up the production management of ”Honda Motor Co.,Ltd. (following Honda)” even of that in the case, and the class to appreciate that importance with ranking of technology in management is taken.

Style⟩ [形態] �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ Finish ” Introduction to new business ” or ” Advanced Lecture in

management Theory of new business ” . Notice⟩ Face a class when a taking lecture person surely does enough preparation

every time. And moreover, be the worst, but one person one time is to do some remarks every time.

Goal⟩

1. The elementary knowledge of management is learned.

2. Ranking of the technology in management and that importance are appreciated.

3. It learns how to solve the various problems which always occur in the spot

of the management through the case study and the business game.

Schedule⟩

1. Guidance

2. Lecture ”What in Management of Technology?”

3. Case report ”Rsearch and development of Honda”

4. Case study ”Plan an overseas factory.” ”A new car development order”

5. Lecture ” Production management and purchasing”

6. Lecture ”Quality control”

7. Case report ”The production management of Honda”,”The supply chain management of Honda”

8. Case study ”Plan an overseas factory.”

9. Case report ”The quality control of Honda”

10. Case study ”Don’t make material defects flare up again.”

11. Lecture ”Marketing strategy”

12. Lecture ”Cross-cultural management”

13. Case report ”Around the products project of Honda - the two-wheeled vehicle ”

14. Case report ”The path of internationalization of Honda”

15. Case report ”The personnel management of Honda”

16. Summary and questions and answers

Evaluation Criteria⟩ It is evaluated in the remark (20%) in class and the middle,

end of a term report (80% in total). Textbook⟩ It doesn’t specially use.The summary and so on is distributed every

time. Reference⟩ Several are introduced in class.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216616

Student⟩ The person who finished ”Introduction to new business” or ” Advanced Lecture in management Theory of new business “ in all the students enrolled in the doctor former term course. And, the number of taking lecture people is restricted to do intensive learning. (20-25 people will be made an upper limit.) It is selected with some means in case of a taking lecture person many.

Contact⟩

⇒ Yamanaka (A410, +81-88-656-7350, yamanaka@ce.tokushima-u.ac.jp)

���� (Office Hour: オフィスアワー:年度ごとに学科の掲示板を参照のこと) ⇒ Morimoto (A417, +81-88-656-7619, emi morimoto@ce.tokushima-u.ac.jp)

���� (Office Hour: 月 ∼ 木 (10 時 ∼ 16 時))

Note⟩ This class invites seven Honda alumnuni who did work in ”research and development”, ”production management”, ”purchasing”, ”quality control”, ”marketing” , ”cross-culutural management”, and so on as an instructor, and it is the class to do case report. Even if it is seen nationwide, the trial done like this is very rare, and it is a precious class. And, an applicant for taking lecture is to attend it in the guidance to hold it in advance.A schedule is indicated separately.